### v2

# Scalable Parameter Design for Superconducting Quantum Circuits with Graph Neural Networks

Abstract
To demonstrate supremacy of quantum computing, increasingly large-scale superconducting quantum computing chips are being designed and fabricated. However, the complexity of simulating quantum systems poses a significant challenge to computer-aided design of quantum chips, especially for large-scale chips. Harnessing the scalability of graph neural networks (GNNs), we here propose a parameter designing algorithm for large-scale superconducting quantum circuits. The algorithm depends on the so-called ’three-stair scaling’ mechanism, which comprises two neural-network models: an evaluator supervisedly trained on small-scale circuits for applying to medium-scale circuits, and a designer unsupervisedly trained on medium-scale circuits for applying to large-scale ones. We demonstrate our algorithm in mitigating quantum crosstalk errors. Frequencies for both single- and two-qubit gates (corresponding to the parameters of nodes and edges) are considered simultaneously. Numerical results indicate that the well-trained designer achieves notable advantages in efficiency, effectiveness, and scalability. For example, for large-scale superconducting quantum circuits consisting of around $870$ qubits, our GNNs-based algorithm achieves $51\%$ of the errors produced by the state-of-the-art algorithm, with a time reduction from $90\ {\rm min}$ to $27\ {\rm sec}$. Overall, a better-performing and more scalable algorithm for designing parameters of superconducting quantum chips is proposed, which initially demonstrates the advantages of applying GNNs in superconducting quantum chips.
Introduction–Quantum computing may have applications in fields including cryptography [1], databases [2], computational chemistry [3], and machine learning [4]. Thus, large-scale quantum information processors are demanded to fully manifest the quantum advantages. Among various quantum computing platforms including neutral atom arrays [5] and trapped irons [6], superconducting quantum computing system based on Josephson junctions is a leading candidate [7, 8, 9, 10, 11, 12]. Qubit numbers and the scales of quantum processors implemented by superconducting quantum electronic circuits (SQECs) are rapidly increasing. The numbers of quantum elements in superconducting quantum computing chip are now more than hundreds and even reach thousands, e.g. in chips of IBM company [13]. Due to the rapid growth of circuit complexity, the analysis and manual design on these large-scale SQECs are becoming challenging. Thus, the electronic design automation for quantum computing systems aided by electronic computers or artificial intelligence is attracting more and more attention [14, 15, 16, 17, 18, 19, 20, 21]. However, most of the discussions are limited to the small or medium scales. Because the dimension of the Hilbert space exponentially scales with the number of qubits in quantum processors [22], the evaluation of the designed SQECs through numerical simulations is becoming a challenging task, which further complicates the automated design towards large-scale quantum processors.
Meanwhile, we notice some neural network models with scalable properties. For example, transformers can deal with texts of different lengths [23], and diffusion models can generate images with different pixels [24]. Within the scalable neural-network models, graph neural networks [25, 26] (GNNs) have demonstrated advantages in a wide range of scientific and engineering applications [27]. By considering qubits as nodes and nearest neighbor coupled-qubit pairs as edges, SQECs can be conceptualized as graphs. This inspires us to achieve automated parameter design for large-scale SQECs by utilizing the scalability of GNNs.
We here focus on the automated design of working parameters in SQECs with given structures. The configuration of parameters on each node and edge can significantly impact the performance of quantum computing. Especially some of the parameters may have graph-dependent effects and need to be configured differently on different nodes and edges. In previous studies, the Snake algorithm was developed for designing parameters in large-scale SQECs [28, 29] and was applied in the context of Google’s Sycamore processor [7]. However, the Snake algorithm ignores the detailed graph structures of the SQECs and relies on multiple calls to traditional optimizers [30]. These characteristics may impact the effectiveness and efficiency of the parameter designing algorithm. In addition to the Snake algorithm, some other algorithms are also proposed for the parameter designing task [31, 32, 33]. However, these other algorithms do not demonstrate their application to large-scale SQECs.
In this study, a GNNs-based algorithm for designing parameters in SQECs is proposed. The realization of the algorithm requires two neural-network models, an evaluator for estimating the quality of parameter assignments and a designer for designing an optimal parameter assignment based on the evaluator’s feedback. The influence of considered parameters is assumed to be localized. Therefore, the evaluator, comprising some graph-based multi-layer perceptrons [34, 35] (MLPs), is supervisedly trained on several small-scale SQECs, e.g., with 6 qubits, to avoid direct simulations on larger quantum processors. Utilizing the errors and corresponding gradients computed by the evaluator, the GNNs-based designer is unsupervisedly trained on randomly generated medium-scale SQECs with around $100$ qubits. Leveraging the scalability of GNNs, the trained designer is directly applied to different graphs in terms of structure or scale. In such way, the parameter designing algorithm can be facilitated for large-scale SQECs containing nearly $1000$ qubits. The proposed algorithm is demonstrated through the task of designing frequency assignments to mitigate quantum crosstalk errors for quantum gates. As quantum crosstalk still exists in ideal systems, it is possible to model the quantum crosstalk errors using numerical simulations [36] instead of real experiments. According to the numerical results, the well-trained GNNs-based designer demonstrates significant advantages over traditional methods in terms of both efficiency and effectiveness for all tested graph scales, ranging from $31$ to $870$ nodes. For instance, in large-scale SQECs comprising $870$ qubits, the trained designer achieves an average crosstalk error of $0.0090$ within $27$ seconds, whereas the Snake algorithm takes $90$ minutes to reach an average crosstalk error of $0.0178$.
The contributions of the paper mainly include the following points. (1) To the best of our knowledge, GNNs are initially applied to the hardware-level design of quantum chips, demonstrating the advantages in performance and scalability. (2) A graph-based scalable evaluator is trained, which can estimate quantum crosstalk errors for large-scale SQECs. While the previous numerical discussions are only limited to small-scale [36]. (3) A GNNs-based designer is trained to assign parameters for SQECs, showcasing the advantages in effectiveness, efficiency, and scalability. The designer performs better than Snake algorithm [28, 29] in both efficiency and effectiveness, while other approaches [31, 32, 33] have not demonstrated their applicability to large-scale SQECs. Note that the focused parameter designing task is a necessary process when utilizing a quantum chip.
Figure 1: Algorithm. In the ’three-stair scaling’ technique, two neural-network models, i.e., the evaluator and the designer, are trained and applied to connect the three scales. The graphic representation of SQECs is shown in a, where graph’s nodes and edges are represented as dots and segments. Throughout the article, nodes or edges are colored with gray when concerning the structure of graphs. While they are colored with rainbow colors when concerning their parameters, where different colors represent different parameters. In b, random parameters are sampled on several different small-scale subgraphs (e.g., the colored parts, each with $6$ qubits) in a larger-scale graph. Quantum crosstalk errors are numerically solved to construct datasets for the evaluator. The evaluator, illustrated in c, utilizes multiple graph-based MLPs to predict quantum crosstalk errors based on the graph structure, parameter assignment, and operation type. The MLPs are supervisedly trained according to the MSE losses and the corresponding gradients. The evaluator trained on small-scale graphs is applied to medium-scale graphs, e.g., d with around $54$ qubits in this study. Then using the evaluated errors and their gradients, the designer shown in e is unsupervisedly trained on random medium-scale graphs such as the graphs in f. The designer assigns parameters for f to obtain d, which integrates GNNs and MLPs to determine optimal node and edge parameters based on the input graphs and trial parameters. Finally, the trained designer can be directly applied to graphs of various structures and scales, including large-scale ones. The example of applying the designer on a large-scale graph is demonstrated with g and h, e.g., a SQEC with $870$ qubits. The comparison of our algorithm (GNNs, with blue markers) and traditional algorithm (Snake, with orange markers) is plotted in i. The tested graphs range from medium-scale to large-scale, and different scales are represented by different markers, e.g., square, triangle and star. More details about the tests can be found in Fig. 2, and the well-trained designer achieves notable advantages in both efficiency and effectiveness.
Algorithm–Our GNNs-based algorithm can design parameters for a coupled-qubit graph conceptualized from a given SQEC, in which various parameters of nodes (qubits) and edges (coupled-qubit pairs) need to be designed for mitigating errors of quantum computing [29, 36]. Some of the parameters are considered graph-independent, which can be designed and optimized individually. While other parameters may have graph-dependent effects, i.e., the parameters of certain node or edge may impact the neighboring nodes and edges. This requires a comprehensive consideration of the entire graph during the designing process. Furthermore, the error dependence on the graph-dependent parameters is complicated, for example, parameters optimized for one node may cause large errors on other nodes. Consequently, powerful algorithms should be implemented to achieve the optimal design of the parameters. Thus, the following studies are based on the graphs conceptualized from SQECs, as schematically shown in Fig. 1(a). From mathematical point of view, the parameter designing task of SQECs can be interpreted as an extension of graph coloring problem [37, 38] (GCP), which involves node and edge colors, continuous coloring, and considerations for the impact of higher-order neighbors.
The GNNs-based parameter designing algorithm considers parameters of nodes and edges, and includes errors from influences of higher-order neighbors (e.g., 2-hop, 3-hop neighbors). As illustrated in Fig. 1, graphs with three scales are involved, i.e., small-scale graphs denoted as stair-$1$, medium-scale graphs denoted as stair-$2$, and large-scale graphs denoted as stair-$3$. The three stairs are connected via two neural-network models: the evaluator and the designer. The evaluator is supervisedly trained on small-scale graphs, and is applied to medium-scale graphs. While the designer is unsupervisedly trained on medium-scale graphs to be applied to large-scale graphs. First, as shown in Fig. 1(b), training data for the evaluator is collected on several small graphs, where a variety of parameter assignments are randomly generated as input labels, and corresponding errors are numerically simulated to serve as output labels. Then the evaluator is trained through supervised learning. As demonstrated in Fig. 1(c), the evaluator employs multiple graph-based MLPs to predict errors based on the graph structure, parameter assignment, and type of quantum operations. When training the evaluator, mean square errors (MSEs) are calculated as loss functions, and the corresponding gradients are used for backpropagation. After training the evaluator, it can be directly applied for parameter assignments on medium-scale graphs, e.g., Fig. 1(d). Unsupervised learning is employed to train the designer using the errors obtained by the evaluator. As illustrated in Fig. 1(e), the designer integrates GNNs and MLPs to determine optimal parameters for nodes and edges based on the input graphs and trial parameters. In each training iteration, a batch of medium-scale graphs is randomly created as input data for the designer, e.g., in Fig. 1(f), some of the nodes and edges are randomly removed from several fundamental graphs. Subsequently, the designer outputs a parameter assignment for each graph as shown with Fig. 1(d), and the corresponding average error is estimated using the trained evaluator. After training the designer on medium-scale random graphs, the scalability of GNNs enables the trained designer to be directly applied to graphs of various structures and scales, including large-scale ones as exemplified with Figs. 1(g) and (h). Additionally, leveraging the parallelism capabilities of machine learning frameworks such as PyTorch [39] and TensorFlow [40], we can further enhance the performance on specific graphs by generating diverse sets of trial parameter assignments and utilizing the designer in batches.
Demonstration–The proposed algorithm is demonstrated with the task of designing frequencies in SQECs to mitigate quantum crosstalk errors. In other words, node and edge frequencies are selected as parameters to be designed by our algorithm. Quantum crosstalk is commonly present in superconducting quantum computing systems [36, 41, 42, 43, 44, 45, 46, 47]. It comes from some fundamental theories in quantum physics, including perturbation theory [48], dressing states [49], and conservation laws, thus crosstalk errors are challenging to be completely eliminated, even in the ideal systems. When numerically simulating quantum crosstalk errors, we consider various factors that may influence the errors, such as residual couplings between high-order neighboring qubits [36, 50], the couplers connected between qubits [51], and higher excitations of these quantum modes [52].
Quantum crosstalk refers to the undesired interference experienced by other qubits when a quantum operation is applied to specific qubits. Three types of quantum operations are involved in the following discussion, including single-qubit $R_{X}$ operation (for nodes), single-qubit $R_{Y}$ operation (for nodes) [53], and two-qubit $R_{XY}$ operation via the $XY$ interaction (for edges) [54]. These three types of quantum operations can be used to build universal quantum computation. Based on the system details outlined in supplementary materials [55], the average crosstalk excitations on a specific node $\mathcal{N}_{k}$ resulting from quantum operations on node $\mathcal{N}_{i}$ (or edge $\mathcal{E}_{ij}$) solely depend on the operation type, graph structure, and frequency assignment. These dependencies delineate the inputs and architecture of the evaluator. Extensive datasets are generated for training, validating, and testing the evaluator. Following the training of the evaluator, the designer undergoes training and testing in accordance with the evaluator’s estimations, showcasing its superiority over other alternative algorithms.
Table 1: The training results of the evaluator.
QOP  [^1] Quantum Operation dataset MSE $r^{2}$
$R_{X}$ training 0.0260 0.9381
validation 0.0268 0.9369
validation (extra) 0.0252 0.9389
testing 0.0280 0.9379
testing (extra) 0.0342 0.9223
$R_{Y}$ training 0.0259 0.9405
validation 0.0286 0.9378
validation (extra) 0.0252 0.9452
testing 0.0291 0.9333
testing (extra) 0.0237 0.9445
$R_{XY}$ training 0.0332 0.9368
validation 0.0301 0.9454
validation (extra) 0.0370 0.9400
testing 0.0261 0.9479
testing (extra) 0.0411 0.9409
Results of the evaluator–After generating the datasets through numerical simulations, the evaluator is trained, validated, and tested in a supervised learning manner. The evaluator consists of multiple MLPs. Each MLP serves for one type of quantum operations and predicts the average crosstalk excitation on a node $\mathcal{N}_{k}$ resulting from operations on node $\mathcal{N}_{i}$ (or edge $\mathcal{E}_{ij}$), where $k\neq i,j$. The inputs of the MLP are the relevant node and edge frequencies denoted by $\vec{\omega}$ and the distances represented by the elements of adjacency matrices, i.e., $A^{p}_{ik}$ (together with $A^{p}_{jk}$ for two-qubit operations). To maintain consistency in the outputs of the MLPs, the crosstalk errors in the simulated datasets are transformed using base-$10$ logarithms. Thus, the losses of the MLPs are determined by the corresponding MSEs, $\mathcal{L}_{O}^{\rm evaluator}=|\Lambda_{O}|^{-1}\sum_{\lambda\in\Lambda_{O}}\delta_{O,\lambda}^{2}$, and
$\delta_{O,\lambda}=|{\rm MLP}_{O}(\vec{\omega}_{\lambda},A^{1}_{\lambda,ik},A^{2}_{\lambda,ik},\cdots)-{\rm log}_{10}(E_{O,\lambda})|,$ (1)
where $O$ is the type of quantum operations, $\lambda$ is a specific index of data in the operation’s dataset $\Lambda_{O}$ with the size $|\Lambda_{O}|$, and $E_{O,\lambda}$ is the simulated crosstalk error. The crosstalk errors smaller than e.g., $1\times 10^{-4}$, or larger than, e.g., $1\times 10^{-1}$, are not concerned, because excessively small crosstalk errors usually do not dominate the error components, while larger ones exhibit strong randomness. Individual MLPs are supervisedly trained for different types of quantum operations, with the resulting MSEs and $r^{2}$ for the training, validation, extra validation, testing and extra testing sets presented in Tab. 1. The coefficient of determination $r^{2}$ exceeds $0.9$ for each case, indicating a strong ability of the evaluator to fit the dataset [56].
Results of the designer–As shown in Fig. 1, during the training of the designer, medium-scale graphs $\mathcal{G}=(\mathcal{N},\mathcal{E})$ with nodes $\mathcal{N}$ and edges $\mathcal{E}$ are randomly generated and input into the designer to output a frequency assignment $\vec{\omega}$ for each graph. The trained evaluator estimates the crosstalk errors for a particular $\vec{\omega}$ on corresponding $\mathcal{G}$. Then the designer’s loss could be calculated,
$\mathcal{L}^{\rm designer}(\mathcal{G},\vec{\omega})$ $=\frac{1}{|\mathcal{N}|}\sum_{\mathcal{N}_{i}\in\mathcal{N}}\sum_{\mathcal{N}_{k}\neq\mathcal{N}_{i}}\hat{E}_{R_{X},\mathcal{N}_{i}\to\mathcal{N}_{k}}$ (2)
$+\frac{1}{|\mathcal{N}|}\sum_{\mathcal{N}_{i}\in\mathcal{N}}\sum_{\mathcal{N}_{k}\neq\mathcal{N}_{i}}\hat{E}_{R_{Y},\mathcal{N}_{i}\to\mathcal{N}_{k}}$
$+\frac{1}{|\mathcal{E}|}\sum_{\mathcal{E}_{i,j}\in\mathcal{E}}\sum_{\mathcal{N}_{k}\neq\mathcal{N}_{i},\mathcal{N}_{j}}\hat{E}_{R_{XY},\mathcal{E}_{ij}\to\mathcal{N}_{k}}.$
First, the estimated crosstalk errors, e.g., $\hat{E}_{R_{X},\mathcal{N}_{i}\to\mathcal{N}_{k}}$, on different nodes $\mathcal{N}_{k}$ are summed to obtain the total crosstalk error of operations on node $\mathcal{N}_{i}$ (or edge $\mathcal{E}_{ij}$). Second, the node(edge)-related errors are subsequently averaged to get the errors for one specific type of quantum operations. Then, the crosstalk errors of all three types of quantum operations are summed up to obtain the comprehensive quantum crosstalk error, serving as the loss function $\mathcal{L}^{\rm designer}(\mathcal{G},\vec{\omega})$ under the specific graph $\mathcal{G}$ and frequency assignment $\vec{\omega}$. Finally, a batch of crosstalk errors for different graphs, together with the parameter assignment on each graph, i.e., $(\mathcal{G},\vec{\omega})$, are averaged to calculate the total loss for unsupervisedly training the designer. In our simulation, only the crosstalk errors up to 4-hop neighbors are taken into account, as the crosstalk errors diminish with distance [55]. The randomly-generated training graphs contain around $54$, $103$ and $218$ nodes, and the training batch size is $288$. Approximately 700,000 training epochs were executed for about $2$ days with an RTX $4090$ GPU. Both node and edge frequencies are encompassed throughout the entire training. Ultimately, the comprehensive crosstalk errors for all three types of quantum operations reach satisfactory levels.
The trained designer can be readily applied to graphs of different structures and scales. The batched utilization strategy, as previously mentioned, can further enhance the designer’s effectiveness. We tested totally $30\times 6$ random graphs with average qubit numbers of $32$, $54$, $103$, $218$, $450$, and $870$, which utilize batch sizes of $512$, $512$, $512$, $64$, $8$, and $2$, respectively. Besides the GNNs-based designer, other methods were also tested for comparisons, including the direct application of a gradient-free optimizer, the direct utilization of gradient descent, and the Snake algorithm. We note that the trained evaluator is also necessary for the traditional methods, because conducting real experiments would require significant time and resources [29], and numerical simulation for SQECs with more than $20$ qubits is also challenging [22]. The effectiveness and efficiency of all the four methods for different graph scales are shown in Fig. 2. All four methods employ the same trained evaluator on the same i9-14900KF CPU. Fig. 2(a) demonstrates that the well-trained GNNs-based designer achieves lower crosstalk errors than all other methods for all the graph scales. Additionally, the GNNs-based designer is exponentially faster than all other methods for all given tasks, as illustrated in Fig. 2(b). Especially for large-scale graphs where direct optimizations are not feasible, the GNNs-based designer achieves slightly more than half the errors of the Snake algorithm in several seconds, while the Snake algorithm would take dozens of minutes.
Figure 2: Effectiveness, efficiency and scalability. These are the detailed versions of Fig. 1(i). a. The average crosstalk errors using different methods on SQECs with various structures and scales. b. The real-world running times of different algorithms. In both a and b, the horizontal axes represent the averaged qubit numbers of the test graphs, ranging from medium to large scales. Here, we test the SQECs with average qubit numbers from $31$ to $870$. The vertical axes correspond to the error and time cost respectively. All axes are expressed in logarithmic coordinates. The application (usage) of our GNNs-based designer is depicted in blue, the Snake algorithm is in orange, the direct application of gradient descent algorithm is in green, and the direct utilization of a gradient-free optimization algorithm is in purple. All the tests are performed using the same CPU and the same evaluator. Only the GNNs-based designer and the Snake algorithm are available for the large-scale graphs. In these figures, the GNNs-based designer demonstrates advantages in terms of effectiveness, efficiency and scalability.
Discussion– This study initially demonstrates the advantages of applying GNNs to the hardware-level optimization of quantum chips, especially for large-scale ones. We apply two graph-based neural-network models named evaluator and designer to tackle the challenge of mitigating quantum crosstalk errors in large-scale SQECs. The evaluator can estimate crosstalk for large-scale SQECs containing thousands of qubits. And the designer outperforms traditional optimization algorithms in efficiency, effectiveness, and scalability. In the future research, additional error sources and variations between qubits may be taken into account, and the evaluator may be trained using experimental data, thus the practicality of the proposed algorithm can be further enhanced.
Acknowledgments–We acknowledge J.-L. Long, Z. Wang, R.-B. Wu, G.-Z. Zhu, Y. Shang, T.-T. Chen, and G.-S. Liu for discussions. This work was support by Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0300201).
Code Availability–The code will be open-soured after the official publication of the paper.
Contents in Supplementary Materials–In Supplementary Materials, we discuss the details of the algorithms and the SQECs based on several references [57, 58, 51, 52, 7, 8, 10, 36, 36, 59, 48, 53, 54, 37, 38, 50, 60, 61, 62, 63, 64, 65, 66, 67, 49, 28, 29, 30, 68], encompassing $8$ pages with $6$ images.